OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2,4) and (5,4).

OpenStudy (jdoe0001):

so, what would be the slope of it?

OpenStudy (anonymous):

how would i find it?

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (jdoe0001):

well \(\large m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

plug in and get the slope

OpenStudy (anonymous):

then the slope would be 0/3???

OpenStudy (anonymous):

i guess yea,

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

0/3 = 0

OpenStudy (jdoe0001):

so the slope is 0 now grab that and plug it in the point-slope form, with either value, say we'll pick (2,4) plug them in, m = slope \(\large y-y_1=m(x-x_1)\)

OpenStudy (jdoe0001):

$$ y-y_1=m(x-x_1)\\ \text{slope of "0" and }(\color{blue}{2}, \color{blue}{4})\\ y-\color{blue}{4}=0(x-\color{blue}{2}) $$

OpenStudy (jdoe0001):

so, what would that give you?

OpenStudy (jdoe0001):

if you solve for "y"

OpenStudy (anonymous):

4....?

OpenStudy (jdoe0001):

y-4 = 0 y = 4 yes, is a flat horizontal line let's grab the 2 points given and graph it 

OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

the equation of the line is y = 4

OpenStudy (jdoe0001):

and is in "slope intercep" form or y = mx+b form y = 4 y = 0x + 4

OpenStudy (anonymous):

Thank you so much!

OpenStudy (jdoe0001):

yw